# L2R: Low-Rank and Lipschitz-Controlled Routing for Mixture-of-Experts

### v1

@@ 3.4 Multi-Anchor Heads
where $\cos\theta_{i,h}=\frac{\bm{q}^{\top}\bm{k}_{i,h}}{\|\bm{q}\|\,\|\bm{k}_{i,h}\|}$. We then aggregate $\{z_{i,h}\}_{h=1}^{H}$ into an expert-level logit $z_{i}$ via log-sum-exp (LSE) pooling (Jia et al., 2012):
@@
Crucially, the bounded SIPS geometry (Section 3.3) makes these prototypes trainable and composable: logits are less dominated by scale, so anchors compete primarily through directional alignment rather than norm inflation, which stabilizes multi-prototype learning and encourages expert cooperation.
@@
Multi-anchor heads only add low-dimensional anchors in $\mathbb{R}^{r}$ and do not replicate the shared projection $\bm{W}_{q}$. As a result, the added parameters scale as $O(NHr)$ and the extra computation is limited to low-rank inner products. A comprehensive efficiency analysis of L2R is provided in Appendix C.
@@
Figure 4: OLMoE training dynamics. Curves show MMLU and HellaSwag accuracies, C4 (Raffel et al., 2020) validation cross-entropy (CE), training CE, and load-balance loss. L2R exhibits clear convergence over 10B tokens and consistently improves MMLU/HellaSwag.
@@ 3.5 Optimization Objective
where $\mathcal{L}_{\text{task}}$ is the task-specific training loss, $\mathcal{L}_{\text{bal}}$ is the standard load-balancing loss (Fedus et al., 2022) that encourages uniform expert usage, and $\mathcal{L}_{z}$ is the router z-loss (Zoph et al., 2022) for stabilizing routing logits during large-scale LLM training. The coefficients $\lambda_{\text{bal}}$ and $\lambda_{z}$ weight the auxiliary terms. Following OLMoE (Muennighoff et al., 2025), we apply $\mathcal{L}_{z}$ for LLM experiments and omit it for vision experiments. Formal definitions of $\mathcal{L}_{\text{bal}}$ and $\mathcal{L}_{z}$ are provided in Appendix D.
@@ 4 Experiments
This section evaluates L2R on both language and vision MoE systems. We first study large-scale pretraining on OLMoE (Muennighoff et al., 2025) and report downstream benchmark performance under the OLMoE evaluation protocol. We then validate generality on a vision MoE system by training a ViT-S (Touvron et al., 2021) backbone on ImageNet-1K (Russakovsky et al., 2015).
@@ 4.1 LLM Experiments: OLMoE Pretraining
L2R is evaluated on OLMoE (Muennighoff et al., 2025), an open MoE language model suite. The backbone and training recipe follow the OLMoE reference configuration with N=64N{=}64 experts and top-k=8k{=}8 sparse activation. All models are trained from scratch for 10B tokens on a distribution-matched subset of the official OLMoE pretraining mixture; subset construction and domain statistics are reported in Appendix E. Unless stated otherwise, L2R uses rank r=2r{=}2 and H=16H{=}16 multi-anchor heads. For SIPS, we use $\gamma{=}1$, $\beta{=}1$ for the query transform (Equation 9) and p=4p{=}4 for the anchor transform (Equation 11). Full hyperparameters and implementation details are provided in Appendix F.1. All runs are conducted on an $8\times$ NVIDIA B200 setup.
@@
The original OLMoE linear router is replaced by each of the compared routing designs. Baselines include the standard Linear router and a cosine scoring method X-MoE (Chi et al., 2022). L2R is evaluated under different scoring geometries, including L2R (Dot), L2R (Cosine), and the proposed L2R (SIPS).
@@
We follow the downstream evaluation protocol and reporting conventions of OLMoE (Muennighoff et al., 2025). Specifically, we track a fixed set of lightweight commonsense and knowledge benchmarks (e.g., MMLU (Hendrycks et al., 2021), ARC-E (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017)) and report Overall as the average score across tasks. This suite is intentionally lightweight for 10B-token training to probe early-stage commonsense and knowledge capabilities while remaining stable and reproducible. Details of task selection and metrics are provided in Appendix H.
@@ 4.1.1 Main Results
Figure 4 summarizes training behavior over 10B tokens. All routers show smooth and monotonic loss reduction, confirming stable convergence under the OLMoE recipe. Throughout training, L2R consistently attains higher MMLU accuracy and remains competitive on HellaSwag. For both C4 (Raffel et al., 2020) validation CE and training CE, L2R closely tracks the angle-sensitive baseline (X-MoE) and converges faster than the Linear router, supporting the benefit of stronger angular discrimination for optimization and generalization. Finally, in terms of load-balance loss, L2R achieves faster and lower convergence, indicating more balanced expert utilization and stable routing dynamics.
Table 1: Overall performance at 10B tokens. Best results are in bold, and second-best results are underlined.
@@
Table 1 reports downstream evaluation results at 10B tokens. L2R with the proposed SIPS achieves the best overall score ($43.4$) and ranks first or second on most benchmarks. Notably, L2R yields a substantial gain on MMLU, improving from $22.7$ (Linear) and $24.6$ (X-MoE) to $27.6$, indicating stronger early-stage knowledge recall and reasoning on lightweight probes. We also observe an exception on ARC-E, where L2R trails the Linear router; consistent with the head ablation in Table 3, the ARC-E drop becomes more pronounced as the number of heads increases, suggesting that the multi-anchor specialization pattern may require longer training to benefit ARC-style multiple-choice questions fully.
@@ 4.1.2 Ablations Study
(a) Varying rank with H=4H{=}4.
(b) Varying heads with r=2r{=}2.
Figure 5: Training dynamics (train cross-entropy loss) for ablations. Both lower rank and more heads yield faster convergence over 10B tokens.
Ablation on Routing Rank $r$.
@@
Figure 6: Variance of pairwise cosine similarity in routing spaces. We compare the layer-averaged variance of pairwise token cosine similarities in the routing space for Linear, X-MoE (Chi et al., 2022), and L2R-SIPS with different rank settings.
Table 3: Ablation on Multi-Anchor Heads ($H$) with Rank r=2r{=}2: Performance with router parameters. Router params over all MoE layers are reported with the percentage relative to the Linear router (2.097M) in parentheses.
@@
Table 3 varies the head count with r=2r{=}2. Increasing $H$ monotonically improves Overall performance, peaking at H=16H{=}16, while incurring only a small router overhead (6.25% of the Linear router). This supports our motivation that multi-anchor heads increase routing expressiveness by allowing each expert to cover multiple modes in the low-rank space at negligible cost. Moreover, Figure 5(b) indicates faster training convergence with larger $H$. We also observe a consistent drop on ARC-E as $H$ increases (aligning with Table 1), suggesting that this multi-mode specialization pattern may require longer training to fully realize ARC-E gains.
@@ 4.1.3 Routing Geometry and Expert Usage Visualizations
Figure 7: Representation-space visualization at a middle layer (Layer 8). PCA projections of token representations in the raw backbone space ($\bm{x}$) and the low-rank routing space ($\bm{q}$). Points are colored by the top-1 routed expert.
@@
Both the raw and projected routing spaces contain low-magnitude tokens around the origin, where routing is inherently noise-sensitive (cosine-style scoring can be particularly brittle to small angular perturbations). SIPS stabilizes this regime by imposing a bounded and smooth magnitude response, avoiding vanishing scores for low-norm queries, thereby improving stability without discarding magnitude information. Further diagnostics and ablations on $\phi(\cdot)$ are provided in Appendix A.
Figure 8: Expert-usage patterns across layers. From top to bottom: top-1 routing frequency, top-$k$ routing frequency, and importance-based routing weights. Each row is a Transformer layer, and each column is an expert.
@@
Figure 8 summarizes routing patterns across layers. Compared to Linear routing, L2R yields a sharper and more structured top-1 map, indicating clearer expert specialization, while maintaining broadly distributed top-$k$ usage, reflecting sustained cooperation under sparse activation. Importantly, regions with high top-1 frequency do not consistently align with peaks in the top-$k$ and importance maps, suggesting that L2R avoids single-expert domination and instead modulates multi-expert contributions via SIPS-controlled magnitude.
@@ 4.2 Experiments with ViT on ImageNet
We further evaluate L2R on a vision MoE system by training a ViT-S (Touvron et al., 2021) backbone on ImageNet-1K (Russakovsky et al., 2015) from scratch for 300 epochs. We follow a standard ViT-S recipe with batch size 512 and apply MoE to the last 4 Transformer blocks. Unless stated otherwise, we use N=16N{=}16 experts with top-k=2k{=}2 sparse activation, and train with only the load-balancing auxiliary loss ($\lambda_{\mathrm{bal}}{=}0.01$). For L2R, we set rank r=2r{=}2 and multi-anchor heads H=4H{=}4, and reuse the same SIPS hyperparameters as in the LLM setting ($\gamma{=}1$, $\beta{=}1$, p=4p{=}4). All runs are conducted on $4\times$ A100 GPUs. Detailed configurations are summarized in Appendix F.2.
Table 4: Top-1 and Top-5 accuracy on ImageNet-1K with ViT-S backbone. Higher is better ($\uparrow$).
@@
Table 4 reports ImageNet results. L2R achieves the best performance, confirming that L2R with the proposed SIPS transfers effectively to vision MoE training. Notably, X-MoE (Chi et al., 2022) slightly underperforms the Linear router in this setting, whereas L2R variants yield consistent gains, suggesting that improving routing geometry is beneficial beyond the language domain.
Table 5: Ablation on Routing Rank $r$ with $H=4$.

### v2

@@ 2.4 Multi-Anchor Heads
where $\cos\theta_{i,h}=\frac{\bm{q}^{\top}\bm{k}_{i,h}}{\|\bm{q}\|\,\|\bm{k}_{i,h}\|}$. We then aggregate $\{z_{i,h}\}_{h=1}^{H}$ into an expert-level logit $z_{i}$ via log-sum-exp (LSE) pooling Jia et al. (2012):
@@
Multi-anchor heads only add low-dimensional anchors in $\mathbb{R}^{r}$ and do not replicate the shared projection $\bm{W}_{q}$. As a result, the added parameters scale as $O(NHr)$ and the extra computation is limited to low-rank inner products. A comprehensive efficiency analysis of L2R is provided in Appendix F.
@@
Figure 4: OLMoE training dynamics. Curves show MMLU and HellaSwag accuracies, C4 Raffel et al. (2020) validation cross-entropy (CE), training CE, and load-balance loss. L2R exhibits clear convergence over 10B tokens and consistently improves MMLU/HellaSwag.
@@ 2.5 Optimization Objective
where $\mathcal{L}_{\text{task}}$ is the task-specific training loss, $\mathcal{L}_{\text{bal}}$ is the standard load-balancing loss Fedus et al. (2022) that encourages uniform expert usage, and $\mathcal{L}_{z}$ is the router z-loss Zoph et al. (2022) for stabilizing routing logits during large-scale LLM training. The coefficients $\lambda_{\text{bal}}$ and $\lambda_{z}$ weight the auxiliary terms. Following OLMoE Muennighoff et al. (2025), we apply $\mathcal{L}_{z}$ for LLM experiments and omit it for vision experiments. Formal definitions of $\mathcal{L}_{\text{bal}}$ and $\mathcal{L}_{z}$ are provided in Appendix G.
@@ 3 Experiments
This section evaluates L2R on both language and vision MoE systems. We first study OLMoE-based language MoE training Muennighoff et al. (2025) and report downstream benchmark performance under the OLMoE evaluation protocol. We then validate generality on a vision MoE system by training a ViT-S Touvron et al. (2021) backbone on ImageNet-1K Russakovsky et al. (2015).
@@ 3.1 LLM Experiments: OLMoE Pretraining
L2R is evaluated on OLMoE Muennighoff et al. (2025), an open MoE language model suite. The backbone and training recipe follow the OLMoE reference configuration with N=64N{=}64 experts and top-k=8k{=}8 sparse activation. All models are trained from scratch for 10B tokens on a distribution-matched subset of the official OLMoE pretraining mixture; subset construction and domain statistics are reported in Appendix H. Unless stated otherwise, L2R uses rank r=2r{=}2 and H=16H{=}16 multi-anchor heads. For SIPS, we use $\gamma{=}1$, $\beta{=}1$ for the query transform (Equation 9) and p=4p{=}4 for the anchor transform (Equation 11). Full hyperparameters and implementation details are provided in Appendix I.1. All runs are conducted on an $8\times$ NVIDIA B200 setup.
@@
The original OLMoE linear router is replaced by each of the compared routing designs. Baselines include the standard Linear router and a cosine scoring method X-MoE Chi et al. (2022). L2R is evaluated under different scoring geometries, including L2R (Dot), L2R (Cosine), and the proposed L2R (SIPS).
@@
We follow the downstream evaluation protocol and reporting conventions of OLMoE Muennighoff et al. (2025). Specifically, we track a fixed set of lightweight commonsense and knowledge benchmarks (e.g., MMLU Hendrycks et al. (2021), ARC-E Clark et al. (2018), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), PIQA Bisk et al. (2020), SciQ Welbl et al. (2017)) and report Overall as the average score across tasks. This suite is intentionally lightweight for 10B-token training to probe early-stage commonsense and knowledge capabilities while remaining stable and reproducible. Details of task selection and metrics are provided in Appendix K.
(a) Varying rank with H=4H{=}4.
(b) Varying heads with r=2r{=}2.
Figure 5: Training dynamics for ablations. Both lower rank and more heads yield faster convergence.
Figure 6: Layer-averaged variance of pairwise token cosine similarities in the routing space.
@@ 3.1.1 Main Results
Figure 4 summarizes training behavior over 10B tokens. All routers show smooth and monotonic loss reduction, confirming stable convergence under the OLMoE recipe. Throughout training, L2R consistently attains higher MMLU accuracy and remains competitive on HellaSwag. For both C4 Raffel et al. (2020) validation CE and training CE, L2R closely tracks the angle-sensitive baseline (X-MoE) and converges faster than the Linear router, supporting the benefit of stronger angular discrimination for optimization and generalization. Finally, in terms of load-balance loss, L2R achieves faster and lower convergence, indicating more balanced expert utilization and stable routing dynamics.
Table 1: Overall performance at 10B tokens.
@@
Note: Best results are in bold, and second-best results are underlined.
@@
Table 1 reports downstream evaluation results at 10B tokens. To isolate the effect of SIPS, we apply SIPS directly to the full-rank Linear router without low-rank projection or multi-anchor heads. Linear (SIPS) improves the overall score from $40.5$ to $42.1$, showing that SIPS alone can provide a clear gain over dot-product scoring. L2R with the proposed low-rank routing and SIPS further achieves the best overall score ($43.4$) and ranks first or second on most benchmarks. Notably, L2R yields a substantial gain on MMLU, improving from $22.7$ (Linear) and $24.6$ (X-MoE) to $27.6$, indicating stronger early-stage knowledge recall and reasoning on lightweight probes. We also observe an exception on ARC-E, where L2R trails the Linear router; consistent with the head ablation in Table 3, the ARC-E drop becomes more pronounced as the number of heads increases, suggesting that the multi-anchor specialization pattern may require longer training to benefit ARC-style multiple-choice questions fully.
@@ 3.1.2 Ablations Study
Ablation on Routing Rank $r$.
Table 3: Ablation on multi-anchor heads $H$ with rank r=2r{=}2.
@@
64 196.6K (9.38%) 26.9 46.9 62.1 40.2 62.1 68.8 42.8
Note: Router parameters are reported over all MoE layers, with percentages relative to the Linear router.
@@
Table 3 varies the head count with r=2r{=}2. Increasing $H$ improves overall performance up to a moderate range, with the best result obtained at H=16H{=}16; however, further increasing $H$ to $64$ does not yield additional gains. This suggests a practical head-selection guideline: a moderate number of heads, such as H=8H{=}8 or H=16H{=}16, provides a favorable trade-off between multi-view routing capacity and anchor redundancy. These results support the effectiveness of the multi-anchor routing pattern, which enriches routing flexibility in the low-rank space without incurring prohibitive router overhead.
@@ 3.1.3 Routing Geometry and Expert Usage Visualizations
Figure 7: Representation-space visualization. PCA projections of token representations in the raw backbone space ($\bm{x}$) and the low-rank routing space ($\bm{q}$). Points are colored by the top-1 routed expert.
@@
Both the raw and projected routing spaces contain low-magnitude tokens around the origin, where routing is inherently noise-sensitive (cosine-style scoring can be particularly brittle to small angular perturbations). SIPS stabilizes this regime by imposing a bounded and smooth magnitude response, avoiding vanishing scores for low-norm queries, thereby improving stability without discarding magnitude information. Further diagnostics and ablations on $\phi(\cdot)$ are provided in Appendix D.
Figure 8: Expert-usage patterns. Each row is a Transformer layer, and each column is an expert. From left to right: top-1 frequency, top-$k$ frequency, and importance-based routing weights.
@@
Figure 8 summarizes routing patterns across layers. Compared to Linear routing, L2R yields a sharper and more structured top-1 map, indicating clearer expert specialization, while maintaining broadly distributed top-$k$ usage, reflecting sustained cooperation under sparse activation. Importantly, regions with high top-1 frequency do not consistently align with peaks in the top-$k$ and importance maps, suggesting that L2R avoids single-expert domination and instead modulates multi-expert contributions via SIPS-controlled magnitude. We further provide quantitative routing diagnostics in Appendix B, including margin-based expert-discrimination metrics and perturbation-based robustness measurements.
@@ 3.2 Experiments with ViT on ImageNet
Table 4: Top-1 and Top-5 accuracy on ImageNet-1K with ViT-S backbone.
@@
We further evaluate L2R on a vision MoE system by training a ViT-S Touvron et al. (2021) backbone on ImageNet-1K Russakovsky et al. (2015) from scratch for 300 epochs. We follow a standard ViT-S recipe with batch size 512 and apply MoE to the last 4 Transformer blocks. Unless stated otherwise, we use N=16N{=}16 experts with top-k=2k{=}2 sparse activation, and train with only the load-balancing auxiliary loss ($\lambda_{\mathrm{bal}}{=}0.01$). For L2R, we set rank r=2r{=}2 and multi-anchor heads H=4H{=}4, and reuse the same SIPS hyperparameters as in the LLM setting ($\gamma{=}1$, $\beta{=}1$, p=4p{=}4). All runs are conducted on $4\times$ A100 GPUs. Detailed configurations are summarized in Appendix I.2.
@@
Table 4 reports ImageNet results. L2R achieves the best performance, confirming that L2R with the proposed SIPS transfers effectively to vision MoE training. Notably, X-MoE Chi et al. (2022) slightly underperforms the Linear router in this setting, whereas L2R variants yield consistent gains, suggesting that improving routing geometry is beneficial beyond the language domain. Additional ablations on routing rank and multi-anchor heads are provided in Appendix C.
4 Conclusion
We introduced Low-rank & Lipschitz-controlled Routing (L2R), a routing framework that improves MoE specialization by jointly redesigning the routing space and scoring geometry. L2R replaces high-dimensional linear routing with a shared low-rank latent space, adopts Saturated Inner-Product Scoring (SIPS) to bound magnitude sensitivity while preserving controlled magnitude for adaptive expert weighting, and extends expressiveness with lightweight multi-anchor heads. Experiments on OLMoE pretraining and ViT-based ImageNet training show consistent gains on most benchmarks, faster convergence, and balanced expert usage, indicating that L2R provides a practical and stable routing recipe across both language and vision MoE systems. We discuss limitations and future directions in Appendix L, and broader impacts in Appendix M.
Appendix A Related Work
Router Design in Sparse MoE.
Most modern sparse MoE systems still rely on raw-space linear routing, where expert logits are produced by a single linear projection in the backbone representation space, e.g., in OLMoE [19], Mixtral [17], Qwen2-MoE [31], and DeepSeek-MoE [8], as well as vision MoE variants such as V-MoE [23], Mod-Squad [4], and ASE [32]. Recent studies revisit router design from multiple angles: improving routing expressiveness (e.g., Router Upcycling [22] and attention-based routing in Yuan 2.0-M32 [29]), and improving stability and utilization via auxiliary objectives and routing variants [26, 18, 11]. Despite these advances, the routing score is still typically produced by linear scoring in a high-dimensional space, which leaves the scoring geometry sensitive to scale and high-dimensional concentration.
Geometry-Aware Latent-Space Routing.
Another line of work performs routing in a latent space and adopts angle-aware similarities [5, 20, 30]. X-MoE [5] projects representations to a lower-dimensional space and uses cosine scoring to alleviate collapse in high-dimensional gating. Different from prior works, our study emphasizes two structural limitations that jointly weaken discriminability under raw-space routing: representation mismatch and angular concentration. We further observe that common rank heuristics (e.g., scaling rank with the number of experts) can still yield an angle-concentrated routing geometry in practice, motivating an ultra-low yet consistent rank across settings. Within this latent-space routing paradigm, L2R introduces SIPS, a Lipschitz-controlled scoring geometry that retains magnitude cues while explicitly bounding their effective influence (instead of discarding magnitude as cosine scoring), and improves expressiveness via lightweight multi-anchor heads.
Appendix B Quantitative Diagnostics of Routing Behavior
B.1 Quantitative Metrics for Expert Discrimination
Table 5: Direct expert-discrimination metrics.
Model Mean Margin $\uparrow$ Low-Margin Rate $\downarrow$
X-MoE 0.0712 0.9424
L2R (Dot)-H1 0.0105 0.9989
L2R (SIPS)-H1 0.1806 0.7036
L2R (SIPS)-H16 0.1718 0.6701
To complement PCA visualizations and routing heatmaps, we introduce two quantitative metrics for expert discrimination. Mean Routing Margin measures the average logit gap between the Top-1 expert and the Top-2 expert, while Low-Margin Rate measures the proportion of tokens whose routing margin is smaller than $0.2$. A larger margin indicates clearer separation between the selected expert and its strongest competitor, whereas a lower low-margin rate indicates fewer tokens near the routing decision boundary. As shown in Table 5, X-MoE and L2R (Dot)-H1 exhibit small margins and high low-margin rates, suggesting weak expert separation. In contrast, L2R with SIPS substantially increases the routing margin and reduces the low-margin rate, providing direct quantitative evidence that SIPS leads to more separated and decisive routing decisions.
B.2 Routing Robustness under Perturbations
To directly validate the stability benefit of SIPS, we apply Gaussian noise with $\sigma=0.02$ to router inputs and measure two perturbation-based routing robustness metrics: Stability of the top-1 routed expert and Top-$k$ Overlap (Jaccard similarity of the selected top-8 expert set).
Table 6: Routing robustness under Gaussian perturbations.
Model Stability $\uparrow$ Top-$k$ Overlap $\uparrow$
Linear 0.9744 0.9679
Linear (SIPS) 0.9853 0.9805
X-MoE 0.9566 0.9683
L2R (Dot)-H1 0.6734 0.8311
L2R (SIPS)-H1 0.9847 0.9915
Note: Higher values indicate more stable routing decisions.
As shown in Table 6, these results provide direct empirical support for the stability claim. Comparing Linear to Linear (SIPS) shows that SIPS improves routing robustness even in the full-rank representation space, indicating that the scoring function itself contributes to stable expert selection. More importantly, low-rank routing with unconstrained dot-product scoring, i.e., L2R (Dot)-H1, becomes highly fragile, with stability dropping to 0.6734 and top-$k$ overlap dropping to 0.8311. In contrast, replacing the unbounded dot product with SIPS largely restores robustness, achieving 0.9847 stability and 0.9915 top-$k$ overlap for L2R (SIPS)-H1. These results clarify the component roles: the low-rank projection reshapes the routing geometry, whereas SIPS makes the resulting low-rank routing stable in practice. We therefore emphasize bounded-domain control of radial sensitivity, rather than unconstrained global smoothness.
Appendix C Ablation Study for ImageNet
We provide ImageNet-1K ablation studies to analyze the effects of the routing rank $r$ and the number of multi-anchor heads $H$ in vision MoE training.
Table 7: Ablation on routing rank $r$ with $H=4$ on ImageNet-1K.